# Deriving a parton shower for jet thermalization in QCD plasmas

### v1

@@ I Introduction
In this work, we present a new MC algorithm that exactly reproduces the dynamics of the linearized EKT. It consistently incorporates energy loss, thermalization, Bose/Fermi statistics, the tracking of recoils and holes, and $2!\to!1$ particle merging. This enables a fully microscopic, first-principles description of jet energy loss and thermalization in the QGP. Furthermore, leveraging the multi-particle capabilities of MC simulations, we compute for the first time the two-particle correlation within the parton shower framework, revealing correlations beyond molecular chaos.
@@ II Theoretical framework
Following the EKT framework [38], the evolution of a small perturbation $\delta f(t,{\bm{x}},{\bm{p}})$ in an equilibrium background $n(p)$ is described by the linearized Boltzmann equation [42, 44],
$\left(\partial_{t}+{\bm{v}}\cdot\nabla_{{\bm{x}}}\right)\delta f_{a}(t,{\bm{x}},{\bm{p}})=\delta C_{a}^{2\leftrightarrow 2}+\delta C_{a}^{1\leftrightarrow 2}\,$, (1)
where $\delta C_{a}^{2\leftrightarrow 2}[\{n,\delta f\}](t,{\bm{x}},{\bm{p}})$ and $\delta C_{a}^{1\leftrightarrow 2}[\{n,\delta f\}](t,{\bm{x}},{\bm{p}})$ describe linearized elastic and inelastic collisions, respectively, and $a$ denotes the parton flavor (quark or gluon). The equilibrium distribution for bosons and fermions is $n_{a}(t,{\bm{x}},{\bm{p}})=1/(e^{p/T}\mp 1)$. Inelastic $1\leftrightarrow 2$ processes are given by
@@
where $\pm_{a}$ sign varies for bosons and fermions. Elastic collisions are given by
@@
where $\int{\rm d}\Omega^{2\leftrightarrow 2}=\int_{k,p^{\prime},k^{\prime}}\delta^{4}(p^{\mu}+k^{\mu}-p^{\prime\mu}-k^{\prime\mu})$, with $\int_{p}=\int_{{\bm{p}}}\frac{1}{2E_{p}}$, $\int_{{\bm{p}}}=\int\frac{d^{3}{\bm{p}}}{(2\pi)^{3}}$, and $|\mathcal{M}^{ab}_{cd}|^{2}$ are the HTL regulated leading-order 2-2 matrix elements, and the associated statistical factor is
$\begin{split}&\delta\mathcal{F}^{ab}_{cd}({\bm{p}},{\bm{k}};{\bm{p}}^{\prime},{\bm{k}}^{\prime})=\\ &+\delta f_{c}({\bm{p}}^{\prime})\left[n_{d}(1\pm n_{a})(1\pm n_{b})\mp_{c}n_{a}n_{b}(1\pm n_{d})\right]\\ &+\delta f_{d}({\bm{k}}^{\prime})\left[n_{c}(1\pm n_{a})(1\pm n_{b})\mp_{d}n_{a}n_{b}(1\pm n_{c})\right]\\ &-\delta f_{a}({\bm{p}})\left[n_{b}(1\pm n_{c})(1\pm n_{d})\mp_{a}n_{c}n_{d}(1\pm n_{b})\right]\\ &-\delta f_{b}({\bm{k}})\left[n_{a}(1\pm n_{c})(1\pm n_{d})\mp_{b}n_{c}n_{d}(1\pm n_{a})\right]\,.\end{split}$ (5)
We omit the momentum dependence, which can be inferred from the indices $(a,b,c,d)$ corresponding to $(a,{\bm{p}})$, $(b,{\bm{k}})$, $(c,{\bm{p}}^{\prime})$, $(d,{\bm{k}}^{\prime})$.
In the following, we introduce the parton shower formulation of the linearized Boltzmann equation. In this work, we neglect space-time inhomogeneities. We first introduce the shorthand notation
@@
where the right-hand side can be read off from eqs.˜2, 3, 4 and 5, distinguishing real and virtual momentum exchanges. Virtual collisions do not change momentum and are therefore directly proportional to $\delta f_{a}({\bm{p}})$. We rewrite eq.˜1 as
@@
where we used the initial condition $\delta f_{a0}=\delta f_{a}(t_{0},{\bm{p}})$, and
$\begin{split}&\Delta_{a}(t,{\bm{p}})=\exp\left[-\int^{t}_{t_{0}}{\rm d}t^{\prime}(\delta C^{1\leftrightarrow 2,{\rm v}}_{a}+\delta C^{2\leftrightarrow 2,{\rm v}}_{a})\right]\,.\end{split}$ (8)
Equation˜7 is the parton shower formulation of the linearized Boltzmann equation, where $\Delta_{a}(t,{\bm{p}})$ is the no-collision probability, analogous to the Sudakov factor in vacuum jet showers [55]. The no-collision probability and the real-collision terms $\delta C^{1\leftrightarrow 2,{\rm r}}_{a}$, $\delta C^{2\leftrightarrow 2,{\rm r}}_{a}$ include splitting, merging, recoil, and hole particles, and therefore go beyond typical MC implementations of elastic [50, 51, 56, 57] and inelastic processes [58, 59, 60, 17]. The appearance of holes (negative particles) becomes more transparent here, as $\delta C^{,{\rm r}}$ contains negative terms. Solving eq.˜7 is equivalent to solving the original linearized Boltzmann equation and therefore provides a consistent description of the equilibration of perturbations in QCD plasmas. When eq.˜7 is solved using MC techniques, initial particles are sampled from $\delta f_{a0}$ and subsequently undergo real collision processes according to the no-emission probability and real-collision kernels.
@@ III High-energy limit
$\begin{split}\delta C^{1\leftrightarrow 2}&\approx\int_{0}^{1}{\rm d}z\left[\tfrac{1}{z^{3}}\Gamma(z,\tfrac{{\bm{p}}}{z})\delta f(t,\tfrac{{\bm{p}}}{z})\right.\\ &-\left.\tfrac{1}{2}\Gamma(z,{\bm{p}})\delta f(t,{\bm{p}})\right]\,.\end{split}$ (9)
We integrate out the angle and introduce the energy density $D(t,x;p_{0})=\nu_{g}\int_{{\bm{p}}}\tfrac{|{\bm{p}}|}{p_{0}}\delta(\tfrac{|{\bm{p}}|}{p_{0}}-x)\delta f(t,{\bm{p}})$, arriving at the well-known turbulent cascade evolution [61, 62, 63, 42]
@@
where the no-emission probability is $\Delta(t,{\bm{p}})=\exp[-\int_{t_{0}}^{t}{\rm d}t^{\prime}\int_{0}^{1}{\rm d}z\tfrac{1}{2}\Gamma(z,{\bm{p}})]$. For a single parton initial condition, $D_{0}=x\delta(1-x)$, the first few splittings are
$\begin{split}&D(t,x;p_{0})=\Delta(t,x{\bm{p}}_{0})x\delta(1-x)\\ &\quad+\int_{t_{0}}^{t}{\rm d}t_{1}\Delta(t-t_{1},x{\bm{p}}_{0})x\Gamma(x,{\bm{p}}_{0})\Delta(t_{1},{\bm{p}}_{0})+\dots\end{split}$ (12)
The first line represents the case with no emissions throughout the evolution, while the second line describes exactly one emission at $t_{1}$, with a splitting fraction $x$, where the two Sudakov factors prevents further emissions before and after. In the parton shower algorithm, first, the splitting time is determined by sampling the Sudakov $\Delta(t_{\rm next},{\bm{p}})$. In case of splitting, $t_{\rm next}<L$, two new particles are created (note the factor of 2 difference between the Sudakov’s exponent and splitting rate in eq.˜11) with zp,(1z)pzp,(1-z)p energies, where $z$ is sampled from the rate $\tfrac{1}{2}\Gamma(z,p)$. This procedure is repeated for all new particles until $t_{\rm next}>L$. A sampling of a final Sudakov ensures that particles do not split after their last creation, $\Delta(L-t_{\rm last},p^{\prime})$. Energy-momentum conservation is encoded in the collinear momentum reconstruction, while the particle number is not conserved. In practice, the MC implements the evolution of the number density $D(x)/x$. This algorithm can be cross-checked against the time evaluation in eq.˜12, and the numerical evolution of eq.˜10.
@@
We evaluated eq.˜10 using a differential equation solver and eq.˜11 with a parton shower. We use the high-energy splitting rate ($T\ll p$), $\Gamma(z,{\bm{p}})=\frac{\alpha_{s}}{2\pi}\frac{C_{A}}{z(1-z)}\sqrt{\frac{C_{A}\hat{q}}{z(1-z)p}}$, where $\hat{q}=\frac{g^{2}C_{A}Tm_{D}^{2}}{2\pi}$, and a single 100 GeV gluon as the initial condition. This splitting rate is the usual choice of medium cascades, and it reproduces the EKT rates in the deep LPM limit (up to logarithmic corrections to $\hat{q}$). This splitting rate has an infrared pole, and thus the separate terms in eqs.˜10 and 11 are divergent. The overall sum is, however, finite as divergences cancel in gain and loss terms. Therefore, we introduced an infrared regulator for both algorithms such that every splitting involves particles with energies bigger than $E_{\text{min}}$. [^2] Common Boltzmann equation solvers implicitly regulate infrared divergences through discrete and finite momentum grids. The number density is evaluated through the histogram
@@
Figure˜1 shows our results with the two different algorithms for different propagation lengths, gray for the traditional diffeq. solver, and color for the new MC. Numerical uncertainties are present for the longest propagation (oscillating gray points). For short paths, most of the energy is in the initial perturbation, while later the cascade quickly transfers energy to the lowest modes. At late time the initial peak disappears and soft modes pile up at the infrared cutoff. Since in the high-energy limit we neglected merging processes, detailed balance is broken and the gluon distribution does not thermalize. In the next section, we derive the full EKT evolution, which includes both splittings and mergings, recovering thermalization at late times.
@@ IV Thermalization of medium induced shower
    which allows both gain and loss contributions (positive and negative terms). The loss contribution is accounted for by introducing hole gluons that propagate like ordinary ones but with 1-1 weights. Hole-particles splitting generates additional hole particles, and the merging of a hole particle produces a normal particle.
The splitting kernels have infrared poles, which are regulated by introducing an energy cutoff, requiring the energy of all splitting legs to be larger than $E_{\rm min}$.
@@
Figure˜2 compares the numerical evaluation of eq.˜7 using the EKT differential-equation solver with the new MC parton shower. Both approaches yield identical distributions. For clarity, we employed the same high-energy splitting rates and neglected elastic collisions, leaving their numerical implementation for future work. We recover a thermal distribution near $p\!\sim\!T$ characterizing the equilibration of the perturbation [42]. Excessive splittings and mergings at low momenta lead to a large number of low-energy particles, slowing down the simulation. The relatively large infrared cutoff $E_{\text{min}}\!=\!1$ GeV avoids this issue, while still allowing for a clear thermal peak. This value is lower than the usual cutoff, $\approx\!5$ GeV, used in other works. Such cutoffs are common in phenomenological applications, as below $\sim\!1$ GeV, non-perturbative and hadronization effects become important. For reference, the asymptotic late-time distribution, $\delta f_{\rm eq}({\bm{p}},{\bm{x}})=\delta T\partial_{T}n_{B}(p)$, is also shown with a dotted line, where $\frac{p_{0}}{V}=\frac{4\pi^{2}}{30}T^{3}\delta T$. There are clear deviations from equilibrium at intermediate times illustrating the need for out-of-equilibrium kinetic evolution instead of instantaneous thermalization and hydro wakes in phenomenological applications.

### v2

@@ I Introduction
In this work, we present a new MC algorithm that exactly reproduces the dynamics of the linearized EKT. It consistently incorporates energy loss, thermalization, Bose/Fermi statistics, the tracking of recoils and holes, and $2\!\to\!1$ particle merging. This enables a fully microscopic, first-principles description of jet energy loss and thermalization in the QGP. Furthermore, leveraging the multi-particle capabilities of MC simulations, we compute for the first time the two-particle correlation within the parton shower framework, revealing correlations beyond molecular chaos.
@@ II Theoretical framework
Following the EKT framework [6], the evolution of a small perturbation $\delta f(t,{\boldsymbol{x}},{\boldsymbol{p}})$ in an equilibrium background $n(p)$ is described by the linearized Boltzmann equation [54, 67],
$\left(\partial_{t}+{\boldsymbol{v}}\cdot\nabla_{{\boldsymbol{x}}}\right)\delta f_{a}(t,{\boldsymbol{x}},{\boldsymbol{p}})=\delta C_{a}^{2\leftrightarrow 2}+\delta C_{a}^{1\leftrightarrow 2}\,$, (1)
where ${\boldsymbol{v}}\!=\!{\boldsymbol{p}}/p$, $\delta C_{a}^{2\leftrightarrow 2}[\{n,\delta f\}](t,{\boldsymbol{x}},{\boldsymbol{p}})$ and $\delta C_{a}^{1\leftrightarrow 2}[\{n,\delta f\}](t,{\boldsymbol{x}},{\boldsymbol{p}})$ describe linearized elastic and inelastic collisions, respectively, and $a$ denotes the parton flavor (quark or gluon). The equilibrium distribution for bosons and fermions is $n_{a}(t,{\boldsymbol{x}},{\boldsymbol{p}})=1/(e^{p/T}\mp 1)$. Inelastic $1\leftrightarrow 2$ processes are given by
@@
where $\pm_{a}$ sign varies for bosons and fermions and the $\boldsymbol{x}$} dependence is local and implicit. Elastic collisions are given by
@@
where $\int{\rm d}\Omega^{2\leftrightarrow 2}=\int_{k,p^{\prime},k^{\prime}}\delta^{4}(p^{\mu}+k^{\mu}-p^{\prime\mu}-k^{\prime\mu})$, with $\int_{p}=\int_{{\boldsymbol{p}}}\frac{1}{2E_{p}}$, $\int_{{\boldsymbol{p}}}=\int\frac{d^{3}{\boldsymbol{p}}}{(2\pi)^{3}}$, and $|\mathcal{M}^{ab}_{cd}|^{2}$ are the HTL regulated leading-order 2-2 matrix elements, and the associated statistical factor is
$\begin{split}&\delta\mathcal{F}^{ab}_{cd}({\boldsymbol{p}},{\boldsymbol{k}};{\boldsymbol{p}}^{\prime},{\boldsymbol{k}}^{\prime})=\\ &+\delta f_{c}({\boldsymbol{p}}^{\prime})\left[n_{d}(1\pm n_{a})(1\pm n_{b})\mp_{c}n_{a}n_{b}(1\pm n_{d})\right]\\ &+\delta f_{d}({\boldsymbol{k}}^{\prime})\left[n_{c}(1\pm n_{a})(1\pm n_{b})\mp_{d}n_{a}n_{b}(1\pm n_{c})\right]\\ &-\delta f_{a}({\boldsymbol{p}})\left[n_{b}(1\pm n_{c})(1\pm n_{d})\mp_{a}n_{c}n_{d}(1\pm n_{b})\right]\\ &-\delta f_{b}({\boldsymbol{k}})\left[n_{a}(1\pm n_{c})(1\pm n_{d})\mp_{b}n_{c}n_{d}(1\pm n_{a})\right]\,.\end{split}$ (5)
We omit the momentum dependence, which can be inferred from the indices $(a,b,c,d)$ corresponding to $(a,{\boldsymbol{p}})$, $(b,{\boldsymbol{k}})$, $(c,{\boldsymbol{p}}^{\prime})$, $(d,{\boldsymbol{k}}^{\prime})$, while the $\boldsymbol{x}$} dependence is local an implicit.
In the following, we introduce the parton shower formulation of the linearized Boltzmann equation. In this work, we neglect space-time inhomogeneities, and so $\delta f(t,{\boldsymbol{p}})=\int{\rm d}^{3}{\boldsymbol{x}}\,\delta f(t,{\boldsymbol{x}},{\boldsymbol{p}})$. We first introduce the shorthand notation
@@
where the right-hand side can be read off from eqs.˜2, 3, 4 and 5, distinguishing real and virtual momentum exchanges. Virtual collisions do not change momentum and are therefore directly proportional to $\delta f_{a}({\boldsymbol{p}})$. We rewrite eq.˜1 as
@@
where we used the initial condition $\delta f_{a0}=\delta f_{a}(t_{0},{\boldsymbol{p}})$, and
$\begin{split}&\Delta_{a}(t,{\boldsymbol{p}})=\exp\left[-\int^{t}_{t_{0}}{\rm d}t^{\prime}(\delta C^{1\leftrightarrow 2,{\rm v}}_{a}+\delta C^{2\leftrightarrow 2,{\rm v}}_{a})\right]\,.\end{split}$ (8)
Equation˜7 is the parton shower formulation of the linearized Boltzmann equation, where $\Delta_{a}(t,{\boldsymbol{p}})$ is the no-collision probability, analogous to the Sudakov factor in vacuum jet showers [31]. The no-collision probability and the real-collision terms $\delta C^{1\leftrightarrow 2,{\rm r}}_{a}$, $\delta C^{2\leftrightarrow 2,{\rm r}}_{a}$ include splitting, merging, recoil, and hole particles, and therefore go beyond typical MC implementations of elastic [64, 7, 40, 29] and inelastic processes [65, 60, 24, 15]. The appearance of holes (negative particles) becomes more transparent here, as $\delta C^{,{\rm r}}$ contains negative terms. Solving eq.˜7 is equivalent to solving the original linearized Boltzmann equation and therefore provides a consistent description of the equilibration of perturbations in QCD plasmas. When eq.˜7 is solved using MC techniques, initial particles are sampled from $\delta f_{a0}$ and subsequently undergo real collision processes according to the no-emission probability and real-collision kernels.
@@ III High-energy limit
$\begin{split}\delta C^{1\leftrightarrow 2}&\approx\int_{0}^{1}{\rm d}z\left[\tfrac{1}{z^{3}}\Gamma(z,\tfrac{{\boldsymbol{p}}}{z})\delta f(t,\tfrac{{\boldsymbol{p}}}{z})\right.\\ &-\left.\tfrac{1}{2}\Gamma(z,{\boldsymbol{p}})\delta f(t,{\boldsymbol{p}})\right]\,.\end{split}$ (9)
We integrate out the angle and introduce the energy density $D(t,x;p_{0})=\nu_{g}\int_{{\boldsymbol{p}}}\tfrac{|{\boldsymbol{p}}|}{p_{0}}\delta(\tfrac{|{\boldsymbol{p}}|}{p_{0}}-x)\delta f(t,{\boldsymbol{p}})$, arriving at the well-known turbulent cascade evolution [13, 14, 47, 54]
@@
where the no-emission probability is $\Delta(t,{\boldsymbol{p}})=\exp[-\int_{t_{0}}^{t}{\rm d}t^{\prime}\int_{0}^{1}{\rm d}z\tfrac{1}{2}\Gamma(z,{\boldsymbol{p}})]$. For a single parton initial condition, $D_{0}=x\delta(1-x)$, the first few splittings are
$\begin{split}&D(t,x;p_{0})=\Delta(t,x{\boldsymbol{p}}_{0})x\delta(1-x)\\ &\quad+\int_{t_{0}}^{t}{\rm d}t_{1}\Delta(t-t_{1},x{\boldsymbol{p}}_{0})x\Gamma(x,{\boldsymbol{p}}_{0})\Delta(t_{1},{\boldsymbol{p}}_{0})+\dots\end{split}$ (12)
The first line represents the case with no emissions throughout the evolution, while the second line describes exactly one emission at $t_{1}$, with a splitting fraction $x$, where the two Sudakov factors prevents further emissions before and after. In the parton shower algorithm, first, the splitting time is determined by sampling the Sudakov $\Delta(t_{\rm next},{\boldsymbol{p}})$. In case of splitting, $t_{\rm next}<L$, two new particles are created (note the factor of 2 difference between the Sudakov’s exponent and splitting rate in eq.˜11) with zp,(1z)pzp,(1-z)p energies, where $z$ is sampled from the rate $\tfrac{1}{2}\Gamma(z,p)$. This procedure is repeated for all new particles until $t_{\rm next}>L$. A sampling of a final Sudakov ensures that particles do not split after their last creation, $\Delta(L-t_{\rm last},p^{\prime})$. Energy-momentum conservation is encoded in the collinear momentum reconstruction, while the particle number is not conserved. In practice, the MC implements the evolution of the number density $D(x)/x$. This algorithm can be cross-checked against the time evaluation in eq.˜12, and the numerical evolution of eq.˜10.
@@
We evaluated eq.˜10 using a differential equation solver and eq.˜11 with a parton shower. We use the high-energy splitting rate ($T\ll p$), $\Gamma(z,{\boldsymbol{p}})=\frac{\alpha_{s}}{2\pi}\frac{C_{A}}{z(1-z)}\sqrt{\frac{C_{A}\hat{q}}{z(1-z)p}}$, where $\hat{q}=\frac{g^{2}C_{A}Tm_{D}^{2}}{2\pi}$, and a single 100 GeV gluon as the initial condition. This splitting rate is the usual choice of medium cascades, and it reproduces the EKT rates in the deep LPM limit (up to logarithmic corrections to $\hat{q}$). This splitting rate has an infrared pole, and thus the separate terms in eqs.˜10 and 11 are divergent. The overall sum is, however, finite as divergences cancel in gain and loss terms. Therefore, we introduced an infrared regulator for both algorithms such that every splitting involves particles with energies bigger than $E_{\text{min}}$. For not too late times, distributions far from $E_{\text{min}}$ are independent of the cutoff. [^2] Common Boltzmann equation solvers implicitly regulate infrared divergences through discrete and finite momentum grids. The number density is evaluated through the histogram
@@
Figure˜1 shows our results with the two different algorithms for different propagation lengths, gray for the traditional diffeq. solver, and color for the new MC. Numerical uncertainties are present for the longest propagation (oscillating gray points). For short paths, most of the energy is in the initial perturbation, while later the cascade quickly transfers energy to the lowest modes. At late time the initial peak disappears and soft modes pile up at the infrared cutoff. Varying $E_{\text{min}}$ changes the late-time position of the peak. Since in the high-energy limit we neglected merging processes, detailed balance is broken and the gluon distribution does not thermalize. In the next section, we derive the full EKT evolution, which includes both splittings and mergings, recovering thermalization at late times.
@@ IV Thermalization of medium induced shower
    which allows both gain and loss contributions (positive and negative terms). The loss contribution is accounted for by introducing a label that distinguishes whether a particle is a hole or not. Hole particles split into hole daughters, while the merging of two holes produces a normal particle. The mixing is also possible by following $g+g\leftrightarrow g$ with their respective signs. When particles are fed to e.g. hadronization, labels are inherited to the offspring. When final particles are analyzed, histogram entries can receive positive or negative contributions depending on whether the particle is a hole.
The splitting kernels have infrared poles, which are regulated by introducing an energy cutoff, requiring the energy of all splitting legs to be larger than $E_{\rm min}$. Other regulator definitions are also possible e.g. regulating $z$. The resulting $\delta f(p)$ should be reliable far from the cutoff $p\gg E_{\text{min}}$. We postpone further studies of infrared sensitivity in EKT to upcoming works.
@@
Figure˜2 compares the numerical evaluation of eq.˜7 using the EKT differential-equation solver with the new MC parton shower. Both approaches yield identical distributions. For clarity, we employed the same high-energy splitting rates and neglected elastic collisions, leaving their numerical implementation for future work. We recover a thermal distribution near $p\!\sim\!T$ characterizing the equilibration of the perturbation [54]. Excessive splittings and mergings at low momenta lead to a large number of low-energy particles, slowing down the simulation. The relatively large infrared cutoff $E_{\text{min}}\!=\!1$ GeV avoids this issue, while still allowing for a clear thermal peak. [^3] Generally speaking, the peak arises from a combination of the thermal bump and the pile-up of particles with $p\sim E_{\text{min}}$. We numerically verified that for $E_{\text{min}}\lesssim T$ only the thermal peak contributes. This value is lower than the usual cutoff, $\approx\!5$ GeV, used in other works. Such cutoffs are common in phenomenological applications, as below $\sim\!1$ GeV, non-perturbative and hadronization effects become important. Differential equation solvers on the other hand use much lower cutoffs i.e. $E_{\text{min}}\ll T$. For reference, the asymptotic late-time distribution, $\delta f_{\rm eq}({\boldsymbol{p}},{\boldsymbol{x}})=\delta T\partial_{T}n_{B}(p)$, is also shown with a dotted line, where $\frac{p_{0}}{V}=\frac{4\pi^{2}}{30}T^{3}\delta T$. There are clear deviations from equilibrium at intermediate times illustrating the need for out-of-equilibrium kinetic evolution instead of instantaneous thermalization and hydro wakes in phenomenological applications.